OpenStudy (anonymous):

Solve the equation for s: -18s + 24 = -9s – 12

OpenStudy (anonymous):

i got this

OpenStudy (anonymous):

We need the s's on one side and the numbers on the other side. Given -18s + 24 = -9s – 12 We can begin by adding 18s to both sides 24 = 9s -12 Do you know the next step?

OpenStudy (anonymous):

i just dont know how to do it ,?

OpenStudy (anonymous):

... im gonna go find something else productive

OpenStudy (anonymous):

then divide

OpenStudy (anonymous):

-18s +9s=-12-24 -9s=-36 s=(-36)/(-9) s=4

OpenStudy (anonymous):

Next 24 = 9s -12 add 12 to both sides 36 = 9s